In a triangle, one angle is 2 times as large as another. The measure of the third angle is 140degrees greater than that of the smallest angle. Find the measure of each angle.
Answer by edjones(8007)

x, 2x, x+140 are the 3 angles
x+2x+x+140=180
4x+140=180
4x+140-140=180-140
4x=40
x=10
2x=20
x+140=150
.
Ed